OpenStudy (anonymous):

solve quadratic equations y=2x-16 y=2^2-3x-12

OpenStudy (stamp):

I would use either the quadratic formula, or set the equations equal to zero and factor to find the solution.

OpenStudy (anonymous):

Is this a system of equations?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, so in the first equation y is solved for. So take that expression (2x-16) and plug it into the second equation where y is Also, is it supposed to be 2x^2 in the second equation?

OpenStudy (anonymous):

y=x^2-3x-12

OpenStudy (anonymous):

Alright, so put 2x-16 where y is 2x-16=x^2-3x-12 can you try to solve for x now?

OpenStudy (anonymous):

Thanks